a cuboidmetal block of dimensions 20*16*12 weight 6 kilogram . find the weight of the block of the same metal of size 10*8*8.
1
Expert's answer
2012-01-10T11:12:50-0500
The density of metal the block consists of is p = M1/V1 = 6/20*16*12 = 1/640, so the weight of the block of the same metal of size 10x8x8 is M2 = p*V2 = (1/640)*(10*8*8) = 1 kg.